News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 8 September 2000 1.00pm EDT

During the last week, Chandra completed the observing schedule as planned.

The observatory passed through the second, third and fourth eclipses of the Fall 2000 season with nominal power and thermal performance. The next 3 eclipses will occur during next week as follows:

Eclipse 5    1:07 pm EDT    10 Sept 2000     79.2 min
Eclipse 6    4:46 am EDT    13 Sept 2000     69.1 min
Eclipse 7    8:29 pm EDT    15 Sept 2000     53.6 min

A real-time ACIS squeegee mode test was conducted at 1am EDT on 5 Sept to test the squeege approach with a back-side CCD (S1). Initial analysis of results shows performance improvement as expected. Other real-time procedures run last week included a routine check of the CPE Warm Start Count (value remained at 1) and the an on-board Ephemeris update.

The planned observations for the coming week are as follows.

Eclipse-5 (79.2 min)                 Sep 10
Radiation Belts
XTE J1550-564           ACIS-S       Sep 11 
NGC 6251                ACIS-I
HR 1442                 HRC-I
M31 (multiple)          HRC-I
AR Lac-1 Eclipse        ACIS-S/HETG
X1543-62                ACIS-S/HETG  Sep 12 
1H 0707-495             ACIS-S/HETG
Eclipse-6 (69.1 min)
Radiation Belts                      Sep 13 
U Gem                   ACIS-S/HETG  Sep 14
AR Lac-3                ACIS-S       Sep 15
Eclipse-7 (53.6 min)
Radiation Belts
AR Lac-2                ACIS-S       Sep 16 
4U 1626-67              ACIS-S/HETG


All spacecraft subsystems continued to operate nominally.

Return to Status Reports